Elorus:

Two functions are shown in the table below: Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 8 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). x = 3 x = 4 x = 5 x = 6

Elorus:

:I

Vocaloid:

To fill out the table, plug each x value into each function, starting with x = 1. For f(x) = −x^2 + 4x + 12, f(1) = -1^2 + 4(1) + 12 = ? For g(x) = x + 8, g(1) = 1 + 8 =? Repeat for all the x values The solution is the x value that makes both functions equal